OpenStudy (anonymous):

A total of $25,000 is invested in two funds paying 5% and 6% annual interest. The combined annual interest is $1,400. How much of the $25,000 is invested into each type of fund?

OpenStudy (anonymous):

You can put this solution on YOUR website! A total of $25,000 is inested in two funds paying 5% and 6% annual interest. The combined annual interest is $1400. How much of the $25,000 is invested in each fund? : let x = amt invested at 6%, then (25000-x) = amt invested at 5% : interest at 6% + interest at 5% = 1400 .06x + .05(25000 - x) = 1400 .06x + 1250 - .05x = 1400 .06x - .05x = 1400 - 1250 .01x = 150 x = 150/.01 x = $15,000 invested at 6% : 25000 - 15000 = $10,000 invested at 5% : Check: .05(10000) + .06(15000) = 1400 .

OpenStudy (anonymous):

give me a medal please